Hi,
Can you help me with this math problem?
Solve the problem.
14.)The sum of three numbers is -5. The first, minus the second, plus 4 times the third, is -20. The third, plus 4 times the first, plus the second, is -20. What are the numbers?
The answer should be (-5,3,-3)

Answer by checkley77(12844)

x+y+z=-5 or y+z=-x-5
x-y+4z=-20
z+4x+y=-20 y+z=-4x-20
-x-5=-4x-20
-x+4x=-20+5
3x=-15
x=-15/3
x=-5 answer.
-5+y+z=-5
y+z=+5-5
y+z=0
y=-z
-5+z+4z=-20
5z=-20+5
5z=-15
z=-15/5
z=-3 answer.
y=-(-3)
y=3 answer.